Question 10
A sample of 10 observations collected in a regression study on three variables, x_1(independent variable), x_2(independent variable and y(dependent variable). The sample resulted in the following data. SSR=27, SST=36.Using a 0.01 level of significance, we conclude that the regression model is significant overall. (Enter 1 if the conclusion is correct. Enter 0 if the conclusion is wrong.)
Answer #1

Regression degrees of freedom = Number of independent variables = 2

Error degrees of freedom = 10 - 2 - 1 = 7

Hence,

F stat = (27/2) / (36/7) = 2.625

Critical F value = 9.547

Hence,

Regression model is not significant overall.

The conclusion is wrong so answer is 0